At a convention there are 7 mathematics instructors, 5 computer science instructors, 3 statistics instructors, and 4 science instructors. If an instructor is selected, find the probability of getting a science instructor or a math instructor



1
Expert's answer
2022-01-24T16:05:45-0500

Total number of people at the convention =19

Probability of getting a science instructor or a maths instructor is.

"=\frac{4}{19}+\frac{7}{19}\\
=\frac{11}{19}\\"